OpenStudy (anonymous):

A compound of nitrogen have largest % composition of nitrogen a:NH3 b:HNO3 c:NH4OH d:(NH2)2CO

OpenStudy (anonymous):

explain me please

OpenStudy (chmvijay):

for that you have to find out molar mass of all these compounds

OpenStudy (chmvijay):

for example NH3 molar mass = 14+3=17 in 17 gram it has 14 gram of N but for 100 gram = 100 *14 /17 =82.35 % like this calculate for all and find in which case % of N is highest

OpenStudy (chmvijay):

hello cant you do this or any problem in understanding this :) @Ryaan

OpenStudy (anonymous):

I solved and got higest% in NH3..is it correct?

OpenStudy (anonymous):

and plus i was little confuse..in finding molecular mass of (NH2)2CO coz this is in bracet..can you help me in this?? means how to calculate

OpenStudy (anonymous):

(NH2)2CO= i calculated this way..14+2=16*2=32..thenC:12+O:16=28..28+32=60 is right?

OpenStudy (chmvijay):

wait let me check

OpenStudy (anonymous):

ok

OpenStudy (chmvijay):

d:(NH2)2CO in this how much % u got

OpenStudy (anonymous):

23.33

OpenStudy (chmvijay):

LOL u should get 46.6 bcoz in this case 60 gram of urea (NH2)2CO contains 28 gram of N right then for 100 gram = 28*60 /100 =46.6

OpenStudy (chmvijay):

there are 2 N atoms in urea molecule (NH2)2 =N2

OpenStudy (chmvijay):

????

OpenStudy (anonymous):

You're right...lol i calculated wrong

OpenStudy (chmvijay):

you are right its NH3 which is having high % of N

OpenStudy (anonymous):

I put 14 instead of 28 then divided nd got:23.3%..which was wrong ..... and thanks for the help..

OpenStudy (chmvijay):

yaaa you now understood where was the really problem good :)yw